A chemist needs to mix a 75% saltwater solution with a 50% saltwater solution to obtain 10 gallons of a 60% saltwater solution. How many gallons of each solution must be used?
Answer by checkley75(3666)

.75x+.5(10-x)=.6*10
.75x+5-.5x=6
.25x=6-5
.25x=1
x=1/.25
x=4 gallons of 60% solution.
10-4=6 gallons of 50% solution.
proof
.75*4+.5*6=6
3+3=6
6=6